Description of Keyboard Mapping

Indexed by Key Sequence

Date: July 24, 1996

Key Sequence Code Sequence in Hex L^{A}TEX
F9  c4  ⋀
F10  ea,eb  ⟹
F11  da  ≡
F12  db  ≠
Ctrl F1  d4  ⊓
Ctrl F2  d5  ⊔
Ctrl F3  d6  \overline{| |}
Ctrl F4  d7  ⨆
Ctrl F5  c5  ⌈
Ctrl F6  c6  ⌉
Ctrl F7  c7  ⌊
Ctrl F8  c8  ⌋
Ctrl F9  c9  (|
Ctrl F10  ca  |)
Ctrl F11  cb  [[
Ctrl F12  cc  ]]
Alt D  a2  Δ
Alt F  a7  Φ
Alt G  a1  Γ
Alt J  a3  Θ
Alt L  a4  Λ
Alt P  a5  Π
Alt Q  a8  Ψ
Alt S  a6  Σ
Alt W  a9  Ω
Alt a  aa  α
Alt b  ab  β
Alt c  bc  χ
Alt d  ad  δ
Alt e  ae  ε
Alt f  bb  φ
Alt g  ac  γ
Alt h  b0  η
Alt j  b1  ϑ
Alt k  b2  κ
Alt l  b3  λ
Alt m  b4  μ
Alt n  b5  ν

Key Sequence Code Sequence in Hex L^{A}TEX
Alt p  b7  π
Alt q  bd  ψ
Alt r  b8  ρ
Alt s  b9  σ
Alt t  ba  τ
Alt w  be  ω
Alt x  b6  ξ
Alt z  af  ζ
Alt F1  e6  ←
Alt F2  e7  −
Alt F3  e8  →
Alt F4  e6,e8  ↔
Alt F5  ec  ↠
Alt F6  ed  ↦
Alt F7  ee  ⇝
Alt F8  ef  ↑
Alt F9  f0  ↓
Alt F10  cd  ∅
Alt F11  ce  ∈
Alt F12  f1  ∉
AltGraph E da  ≡
AltGraph F c4  ⋀
AltGraph G e1  ⪰
AltGraph I  ea,eb  ⟹
AltGraph L df  ⪯
AltGraph M eb  ⇒
AltGraph N f7  ♮
AltGraph P dd  ⊑
AltGraph S  e3  ≃
AltGraph a  c0  ∧
AltGraph b  d8  ⊥
AltGraph e  d9  ≐
AltGraph f  c2  ∀
AltGraph g  e0  ≻
AltGraph i  e7,e8  ⟶
AltGraph l  de  ≺
AltGraph m e8  →
AltGraph n  bf  ¬

Key Sequence Code Sequence in Hex L^{A}TEX
AltGraph o  c1  ∨
AltGraph p  dc  ⊏
AltGraph s  e2  ∼
AltGraph t  c3  ∃
AltGraph u  db  ≠
AltGraph x  f2  ×
AltGraph F1 d0  ∩
AltGraph F2 d1  ∪
AltGraph F3 d2  ⋂
AltGraph F4 d3  ⋃
AltGraph F5 cf  ⊆
AltGraph F6 f9  □
AltGraph F7 fa  ◇
AltGraph F8 fb  ∘
AltGraph F9 fc  •
AltGraph F10 fd  ∥
AltGraph F11 fe  √
AltGraph F12 ff  \begin{matrix}c \\ ◯ \end{matrix}
Shift F1  e9  ⇐
Shift F2  ea  =
Shift F3  eb  ⇒
Shift F4  e9,eb  ⇔
Shift F5  f3  ⊕
Shift F6  f4  ⊖
Shift F7  f5  ⊗
Shift F8  f6  ⊘
Shift F9  f7  ♮
Shift F10  f8  ∞
Shift F11  e4  ≤
Shift F12  e5  ≥